- Press Release
- Dec 1, 2022
NASA Space Station On-Orbit Status 26 February 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Saturday, first weekend rest day for the crew.
Wearing protective garment, CDR/SO Leroy Chiao and FE Salizhan Sharipov completed the regular weekly 3-hour task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]
Sharipov completed another task on the Russian satellite navigation system (ASN-M) in the Service Module (SM), switching and reconnecting cables of the onboard data measuring & storage system (SBI) which he had exchanged on 2/17 for a test of the NPM receiver module units. [To complete the ASN installation for the European ATV (automated transport vehicle), due to be launched in 2006, two additional external ASN antennas will be installed during the Russian EVA-13 on 3/25. For today’s task, the ground first had to turn off the BITS2-12 onboard telemetry measurement system and its VD-SU control mode and later powered it up again. As previously, the work was supported by S-band tagup with ground specialists. ASN-M will use GLONASS satellites (the Russian GPS equivalent) to provide navigational data to the ATV.]
Chiao collected the periodic reading of the cabin air’s current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (carbon dioxide monitoring kit), for calldown (along with the battery status) for use in trending analyses.
Leroy also completed the regular weekly maintenance reboot on the operational PCS (portable computer system) laptops and the bi-monthly restart of the OCA comm router laptop.
Cleaning up from yesterday’s Robotics/SSRMS activities, which completed commissioning of the Canadarm2 for ground control, the CDR disconnected and removed the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab RWS (Robotics Work Station) that supported video camera coverage of the ops.
At ~6:00am EST, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.
At ~9:00am EST, Chiao and Sharipov engaged in their weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and TsUP/Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.
For his “Saturday Science” program today, the Science Officer worked on the SAMS (Space Acceleration Measurement System), using the vacuum cleaner to clean the filter screens of all SAMS drawers and afterwards reactivating the SAMS ICU (interim control unit) in Drawer 1. [The ground then took over with Drawers 1 & 2 RTS (remote triaxial sensor) monitoring of microaccelerations.]
Working off the Russian discretionary task list, the FE performed another session with the ETD-3 experiment (Investigation of the Coordination of Eye and Head Movements) of André Kuipers’ VC6 “Delta” program, before doing any physical exercise. [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing’s plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane. Each step required another prior calibration run, using visual target cues or the calibration unit. The 1-hr. run was supported by tagup with ground specialists.]
As a second task-listed job, Salizhan concluded the latest run of the Molniya-SM/LSO experiment by downloading recorded measurements from the LSO memory to HDD (hard-disk drive) on the EGE-1 laptop for return to Earth. [Objective of Molniya-SM, similar to the French LSO experiment, is to catch and record incidental storm phenomena and other related events in the Earth’s equatorial regions.]
Sharipov conducted the routine daily maintenance of the SOZh environment control & life support system in the SM.
The crew performed their regular daily 2.5-hr. physical exercise program on TVIS, RED exerciser, CEVIS cycle and VELO cycle with bungee cord load trainer. [Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO (today: Day 1 of a new set).]
Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
Another Lesson for Exploration: A paper on ADUM (Advanced Diagnostic Ultrasound in Microgravity) science aboard the ISS in the current (January) issue of the Journal of Trauma concludes that remotely guided FAST (Focused Assessment with Sonography for Trauma) examination by nonphysicians “is possible with excellent clinical results and speed, even with a significantly reduced video frame rate and a 2-second communication latency. A wider application of trauma ultrasound applications for remote medicine on earth appears to be possible and warranted” ( http://www.jtrauma.com)
Weekly Science Update (Expedition Ten — 17th):
GASMAP: Thanks went up to the crew for their successful routine health check of the GASMAP system this week.
Human Research Facility/Workstation (HRF WS): Continuing.
Advanced Ultrasound (ADUM): See comment re “Journal of Trauma” paper, above.
Hand Posture Analyzer (HPA: Nothing new.
In-Space Soldering Investigation (ISSI): Operations are complete.
Pore Formation and Mobility Investigation (PFMI): Nothing new.
Space Acceleration Measurement System (SAMS): Leroy Chiao was thanked for selecting SAMS filter cleaning as this week’s “Saturday Science” activity. The ground deactivated the SAMS drawers so they were ready at his convenience. Also, SAMS’ support has been requested to provide data during the upcoming 16P undocking and 17P docking.
Microgravity Acceleration Measurement System (MAMS): MAMS remains in nominal operations. The payload supported the Progress 16P reboost with data and plots to the Ballistics engineers.
Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.
Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.
Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.
Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.
Renal Stone (RS): Nothing new.
Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.
Foot/Ground Reaction Forces during Space Flight (FOOT): The Science Officer was thanked for taking the time to configure one of the Foot kits last Saturday. “You have successfully assisted us in preparation for Increment 11”.
Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.
Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.
Dust and Aerosol Measurement Feasibility Test (DAFT): The DAFT team sent up its thanks to Dr. Chiao for setting up the P-Trak and the DustTrak to take an air sample near Express Rack 4. Measurements indicated an atmospheric cleanliness on board approaching that of a Clean Room on the ground: “Everything went smoothly during operations. We have looked at the data and video. The video was very helpful which was used it to sync up the times for each instrument with respect to the HEPA filter and sample hose. The readings on both instruments were reading low. They, however, seem credible based on the information we have been told about filter changing on the ISS.”
Biotechnology Specimen Temperature Controller (BSTC): Nothing new.
Yeast Group Activation Packs (Yeast GAP): Nothing new.
Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new.
Earth Science Toward Exploration Research (ESTER): Nothing new.
Serial Network Flow Monitor (SNFM): Nothing new.
Fluid Merging Viscosity Measurement (FMVM): Looking forward to future operations with honey samples.
Space Experiment Module (SEM): Nothing new.
Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.
Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.
Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.
Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.
Miscible Fluids in Microgravity (MFMG): Nothing new.
Educational Payload Operations (EPO): Nothing new.
Capillary Flow Experiment (CFE): Nothing new.
Crew Earth Observations (CEO): Through 2/22 the ground has have received 11,310 of ISS CEO images. An interesting winter view of the northern Chinese industrial city of Shenyang has been posted on NASA’s Earth Observatory website this weekend. Snow cover enhances the compact symmetry of this modern city and contrasts it sharply with the ancient agrarian landscape of the region. The crew continues to acquire many fine images, especially with the long lens. Regretfully, there are few science targets available for them at this time.
Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Vegetation, Galapagos Islands, S. America (Dynamic Event. Weather was predicted to be clear over the Galapagos Islands. Vegetation on the islands typically responds dramatically to strong or weak El Nino events. Last year’s El Nino was of moderate strength. This overpass provided an opportunity to capture the current vegetation conditions on the islands), and Internal waves, Patagonian Shelf (this overpass brought ISS over a complex portion of the Patagonian coastline [small embayments and spits of land]. These coastal features may cause interesting internal wave patterns. Looking to the left of track for the sunglint point).
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- Progress M-51 (16P) undocking — tomorrow, 2/27 (11:08am EST);
- Progress M-52 (17P) launch – Monday, 2/28 (2:09pm EST);
- Progress M-52 (17P) docking – Wednesday, 3/2 (3:15pm EST);
- EVA-13 — 3/25;
- Soyuz TMA-6 (10S) launch — 4/15 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
- Soyuz TMA-5 (9S) undocking — 4/25 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
- LF1 (STS-114) — NET 5/12;
- Progress M-53 (18P) launch — 6/10;
- ULF1.1 (STS-121) — NET 7/10;
- Progress M-54 (19P) launch — 8/24;
- Soyuz TMA-7 (11S) launch — 9/27.
- ISS Orbit (as of this morning, 6:21am EST [= epoch]):
- Mean altitude — 358.1 km
- Apogee height — 361.0 km
- Perigee height — 355.1 km
- Period — 91.70 min.
- Inclination (to Equator) — 51.65 deg
- Eccentricity — 0.0004338
- Solar Beta Angle — 30.9 deg (magnitude increasing)
- Orbits per 24-hr. day — 15.70
- Mean altitude loss in last 24 hours — 75 m
- Revolutions since FGB/Zarya launch (Nov. 98) — 35827
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.